OpenStudy (gabylovesyou):

Daniela teaches Ballet. She charges a one−time processing fee of $3.00 and the cost of each class is shown below. Let x represent the number of classes and y represent the cost of classes. Based on this information, what will it cost, in dollars, to take 6 classes? (Do not put a $ in your answer.) Cost of classes (not including processing fee) Number of Classes, x 1 2 3 4 Cost of Classes, y $16.00 $32.00 $48.00 $64.00

OpenStudy (gabylovesyou):

@ganeshie8

OpenStudy (johnweldon1993):

Well you can see that each class costs 16 dollars... y = 16x where 'x' is the number of classes....so to find the cost for 6 classes...replace 'x' with 6 and at the very end....be sure to add 3 dollars for that processing fee :)

OpenStudy (johnweldon1993):

so actually it should be y = 16x + 3

OpenStudy (gabylovesyou):

do i solve it ?

OpenStudy (johnweldon1993):

yes ma'am

OpenStudy (gabylovesyou):

y = 16x + 3 - 3 y - 3 = 16x/16 y - 3/16 = x

OpenStudy (johnweldon1993):

no, the 'x' would be where you replace with the number of classes.. y = 16x + 3 you want to find out how much it would cost to take 6 classes... so all you would do is replace 'x' with 6 to find the cost 'y' y = 16(6) + 3 what does 'y' = ...?

OpenStudy (gabylovesyou):

99 ?

OpenStudy (johnweldon1993):

perfect...that is how much it would cost to take those lessons for 6 classes